# Some Primal-Dual Theory for Subgradient Methods for Strongly Convex Optimization

### v3

@@ 3.2 Proof of Primal-Dual Convergence Guarantees
where the first inequality uses our inductive result, the second inequality uses the definition of $C_{0}$, and the third inequality bounds the first two summations by noting (i) if $s(x_{k})\neq 0$, then $\delta_{k}(x_{\mathtt{SL}})\geq\tau_{\mathtt{SL}}>0$ and (ii) if $s(x_{k})=0$, then $\delta_{k}(x_{\mathtt{SL}})\geq\inf h_{x_{\mathtt{SL}}}-h_{x_{\mathtt{SL}}}(x_{\mathtt{SL}})$, which may be negative. Rearrangement gives the claim as
$\sum_{k<T:s(x_{k})=0}\lambda_{k}\geq\frac{\tau_{\mathtt{SL}}\sum_{k=0}^{T-1}\lambda_{k}-L_{0}^{2}\sum_{k=0}^{T-1}\lambda_{k}\alpha_{k}-C_{0}}{2(h_{x_{\mathtt{SL}}}(x_{\mathtt{SL}})-\inf h_{x_{\mathtt{SL}}})+\tau_{\mathtt{SL}}}\ .$
@@
$D_{T}+\sum_{k=0}^{T-1}\frac{\lambda_{k}}{2}\mathbb{E}_{\xi}\delta_{k}(x_{\mathtt{OPT}})\leq D_{0}+\sum_{k=0}^{T-1}\frac{L_{0}^{2}\lambda_{k}\alpha_{k}}{2}+\sum_{k=0}^{T-1}\frac{\lambda_{k}}{2}\left(L_{1}\alpha_{k}-1\right)\mathbb{E}_{\xi}\delta_{k}(x_{\mathtt{OPT}})\ .$
@@
$R_{T}(x_{\mathtt{OPT}})+D_{T}+\sum_{k=0}^{T-1}\lambda_{k}\mathbb{E}_{\xi}\delta_{k}(x_{\mathtt{OPT}})\leq\sum_{k=0}^{T-1}L_{0}^{2}\lambda_{k}\alpha_{k}+C_{0}\ .$
Lower bounding each $\mathbb{E}_{\xi}\delta_{k}(x_{\mathtt{OPT}})$ with $s(x_{k})\neq 0$ by zero completes our proof.
@@ 4 Numerical Experiments
In this section, we numerically validate the accuracy of Theorem 3.2 in predicting actual observed performance. Our three main numerical experiments address the impact of varying $\lambda_{k}$, the quality of our new primal-dual stopping criteria, and the accuracy of our $T_{0}$ and $C_{0}$ constants at predicting initial divergences. All of our numerics are implemented in Julia 1.8.5555The source code is available at https://github.com/AshleyLDL/Primal-Dual-Averaging-Coding.
@@
Noting $\|Ax-b\|_{1}$ is $\sum_{i=1}^{n}\|A_{i}\|$-Lipschitz where $A_{i}$ denotes $A$’s $i$th column and $\frac{1}{2}\|Cx-d\|_{2}^{2}$ is $\lambda_{max}(C^{T}C)$-smooth, by (2.9), our Assumptions A-C hold with $L_{0}^{2}=8(\sum_{i=1}^{n}\|A_{i}\|)^{2}$, $L_{1}=4\lambda_{max}(C^{T}C)$ and $\mu=\lambda_{min}(C^{T}C)$. We generate problem instances fixing $m=n=100,x_{0}=0$ and randomly drawing $A,\tilde{C},x_{\mathtt{OPT}}$ with i.i.d. normal entries. To control $\mu$ and $L_{1}$, we set $C=I+\sigma\tilde{C}$ for various selections of $\sigma\geq 0$. When $\sigma=0$, we have $\mu=1$ and $L_{1}=4$. Initially as $\sigma$ increases, $\mu$ decreases while $L_{1}$ increases. To ensure $x_{\mathtt{OPT}}$ is a minimizer and $p_{\star}=0$, we set $b=Ax_{\mathtt{OPT}}$, $d=Cx_{\mathtt{OPT}}$.
@@ 4.3 Accuracy of $C_{0}$ at Predicting Early Iterate Divergence
$\sigma$
0
0.0001
0.001
0.01
0.02
0.05
$L_{1}/\mu$
4
4.022
4.224
6.911
12.107
81.179
$T_{0}$
6
7
7
12
23
161
$C_{0}$
$1.472{\times}10^{5}$
$1.497{\times}10^{5}$
$1.735{\times}10^{5}$
$6.985{\times}10^{5}$
$3.770{\times}10^{6}$
$2.663{\times}10^{23}$
@@
Lastly, we consider settings where the initial iterates diverge rapidly, which our theory addresses via the inclusion of the constant $C_{0}$. Here, we have defined $C=\sigma\tilde{C}+I$, for a randomly Gaussian sampled $\tilde{C}$. As a result, the constants $\mu=\lambda_{min}(C^{T}C)$ and $L_{1}=4\lambda_{max}(C^{T}C)$ depend on $\sigma$. In Table 2, we show the effect $\sigma$ varying from 00 to $0.05$, causing the condition number $L_{1}/\mu$ to grow moderately. As a result, we see $T_{0}$ grow linearly in $L_{1}/\mu$ and $C_{0}$ grows exponentially, exceeding $10^{21}$.
For such problems, our theory predicts the subgradient method with $\alpha_{k}=2/\mu(k+2)$ may diverge in the first $T_{0}$ iterations but should eventually converge at least a $O(1/T)$ rate. Figure 2 numerically confirms this prediction with every performance measure exponentially growing to at least $10^{16}$ as $\sigma$ grows and a decreasing trend beginning before iteration $T_{0}$. We see $\bar{p}_{k}$, $\delta_{k}$ and $R_{k}^{2}(x_{\mathtt{OPT}})$ rapidly converge after $T_{0}$, whereas $p_{k}$ and $d_{k}$ only decrease sublinearly. This slow convergence is likely due to $p_{k}$ and $d_{k}$ being defined as weighted averages, which must slowly dilute the effects of early “bad” iterations. Our theory predicts such exponential divergences can be avoided by ensuring $T_{0}$ (and hence $C_{0}$) are small. For example, setting $\alpha_{0}=1/\mu$ and then $\alpha_{k}=\min\{1/L_{1},2/\mu(k+2)\}$ thereafter rather than $\alpha_{k}=2/\mu(k+2)$ as above ensures $C_{0}=0$. Figure 3 verifies this mitigates the previous diverging behavior.

### v5

@@ 3.2 Proof of Primal-Dual Convergence Guarantees
where the first inequality uses our inductive result, the second inequality uses the definition of $C_{0}$ in (2.10) and that $\delta_{k}(x_{\mathtt{OPT}})\geq 0$, and the third inequality bounds the first two summations as follows: (i) the first sum’s upper bound notes that if $s(x_{k})\neq 0$, then $\delta_{k}(x_{\mathtt{SL}})\geq\tau_{\mathtt{SL}}>0$ and (ii) the second sum’s upper bound notes $L_{1}\alpha_{k}-2\geq-2$ and if $s(x_{k})=0$, then $\delta_{k}(x_{\mathtt{SL}})\geq\inf h_{x_{\mathtt{SL}}}-h_{x_{\mathtt{SL}}}(x_{\mathtt{SL}})$, which may be negative. Rearrangement gives the claim as
$\mathbb{E}_{\xi}\left[\sum_{k<T:s(x_{k})=0}\lambda_{k}\right]\geq\frac{\tau_{\mathtt{SL}}\sum_{k=0}^{T-1}\lambda_{k}-L_{0}^{2}\sum_{k=0}^{T-1}\lambda_{k}\alpha_{k}-C_{0}}{2(h_{x_{\mathtt{SL}}}(x_{\mathtt{SL}})-\inf h_{x_{\mathtt{SL}}})+\tau_{\mathtt{SL}}}\ .$
@@
$D_{T}$ $+\sum_{k=0}^{T-1}\frac{\lambda_{k}}{2}(\mathbb{E}_{\xi}\delta_{k}(x_{\mathtt{OPT}}))+{\color[rgb]{0.0,0.0,0.0}\definecolor[named]{pgfstrokecolor}{rgb}{0.0,0.0,0.0}\pgfsys@color@gray@stroke{0.0}\pgfsys@color@gray@fill{0.0}\sum_{k<T:s(x_{k})>0}\lambda_{k}\mathbb{E}_{\xi}f_{s(x_{k})}(x_{\mathtt{OPT}})}$
$\leq D_{0}+\sum_{k=0}^{T-1}\frac{L_{0}^{2}\lambda_{k}\alpha_{k}}{2}+\sum_{k=0}^{T-1}\frac{\lambda_{k}}{2}\left(L_{1}\alpha_{k}-1\right)\mathbb{E}_{\xi}\delta_{k}(x_{\mathtt{OPT}})\ .$
@@
$R_{T}(x_{\mathtt{OPT}})+D_{T}+\sum_{k=0}^{T-1}\lambda_{k}\mathbb{E}_{\xi}\delta_{k}(x_{\mathtt{OPT}}){\color[rgb]{0.0,0.0,0.0}\definecolor[named]{pgfstrokecolor}{rgb}{0.0,0.0,0.0}\pgfsys@color@gray@stroke{0.0}\pgfsys@color@gray@fill{0.0}+\sum_{k<T:s(x_{k})>0}\lambda_{k}\mathbb{E}_{\xi}}f_{s(x_{k})}(x_{\mathtt{OPT}})\leq\sum_{k=0}^{T-1}L_{0}^{2}\lambda_{k}\alpha_{k}+C_{0}\ .$
Lower bounding each $\mathbb{E}_{\xi}[\delta_{k}(x_{\mathtt{OPT}}){\color[rgb]{0.0,0.0,0.0}\definecolor[named]{pgfstrokecolor}{rgb}{0.0,0.0,0.0}\pgfsys@color@gray@stroke{0.0}\pgfsys@color@gray@fill{0.0}+f_{s(x_{k})}(x_{\mathtt{OPT}})}]$ with $s(x_{k})\neq 0$ by zero and dividing through by $\sum_{k=0}^{T-1}\lambda_{k}$ establishes that
$\frac{1}{\sum_{k=0}^{T-1}\lambda_{k}}\left(R_{T}(x_{\mathtt{OPT}})+D_{T}+\sum_{k<T:s(x_{k})=0}\lambda_{k}\mathbb{E}_{\xi}\delta_{k}(x_{\mathtt{OPT}})\right)\leq\frac{\sum_{k=0}^{T-1}L_{0}^{2}\lambda_{k}\alpha_{k}+C_{0}}{\sum_{k=0}^{T-1}\lambda_{k}}\ .$
Substituting the definitions of $R_{T}$ and $D_{T}$ from (3.7) and (3.8), the left-hand side above is equal to
$\mathbb{E}_{\xi}\left[\frac{\mu}{2}\|x_{T}-x_{\mathtt{OPT}}\|^{2}_{2}+\frac{1}{{\sum_{k=0}^{T-1}\lambda_{k}}}\left(\sum_{k<T:s(x_{k})=0}\lambda_{k}h_{x_{\mathtt{OPT}}}(x_{k})-\inf M^{(T-1)}\right)\right]\ .$
Recalling the primal and dual gaps definitions in (3.3) and (3.4) gives our main theorem.
@@ 4 Numerical Experiments
In this section, we numerically validate the accuracy of Theorem 3.2 in predicting actual observed performance. Our three main numerical experiments address the impact of varying $\lambda_{k}$, the quality of our new primal-dual stopping criteria, and the accuracy of our $T_{0}$ and $C_{0}$ constants at predicting initial divergences. All of our numerics are implemented in Julia 1.8.5666The source code is available at https://github.com/AshleyLDL/Primal-Dual-Averaging-Coding.
@@
Note $\|Ax-b\|_{1}$ is $\|A^{T}\|_{\infty\rightarrow 2}$-Lipschitz [^7] The Lipschitz constant for $\|Ax-b\|_{1}$ follows from the chain rule as its subgradients are combinations of $A$’s rows with weights in $[-1,1]$, so the largest subgradient is $\max_{\|w\|_{\infty}\leq 1}\|A^{T}w\|_{2}=\|A^{T}\|_{\infty\rightarrow 2}$.. However, computing this induced matrix norm is NP-hard [35], so we instead upper bound it by $\sum_{i=1}^{m}\|A_{i}\|$ where $A_{i}$ denotes $A$’s $i$th row. Further noting $\frac{1}{2}\|Cx-d\|_{2}^{2}$ is $\lambda_{max}(C^{T}C)$-smooth, by Lemma 2.4, our Assumptions A-C hold with $L_{0}^{2}=8(\sum_{i=1}^{m}\|A_{i}\|)^{2}$, $L_{1}=4\lambda_{max}(C^{T}C)$ and $\mu=\lambda_{min}(C^{T}C)$. We generate problem instances fixing $m=n=100,x_{0}=0$ and randomly drawing $A,\tilde{C},x_{\mathtt{OPT}}$ with i.i.d. normal entries. To control $\mu$ and $L_{1}$, we set $C=I+\sigma\tilde{C}$ for various selections of $\sigma\geq 0$. When $\sigma=0$, we have $\mu=1$ and $L_{1}=4$. Initially as $\sigma$ increases, $\mu$ decreases while $L_{1}$ increases. To ensure $x_{\mathtt{OPT}}$ is a minimizer and $p_{\star}=0$, we set $b=Ax_{\mathtt{OPT}}$, $d=Cx_{\mathtt{OPT}}$.
@@ 4.3 Accuracy of $C_{0}$ at Predicting Early Iterate Divergence
$\sigma$ 0 0.0001 0.001 0.01 0.02 0.05
$L_{1}/\mu$ 4 4.022 4.224 6.911 12.107 81.179
$T_{0}$ 6 7 7 12 23 161
$C_{0}$ $1.472{\times}10^{5}$ $1.497{\times}10^{5}$ $1.735{\times}10^{5}$ $6.985{\times}10^{5}$ $3.770{\times}10^{6}$ $2.663{\times}10^{23}$
@@
Lastly, we consider settings where the initial iterates diverge rapidly, which our theory addresses via the inclusion of the constant $C_{0}$, defined in (2.10). Here, we have defined $C=\sigma\tilde{C}+I$, for a randomly Gaussian sampled $\tilde{C}$. As a result, the constants $\mu=\lambda_{min}(C^{T}C)$ and $L_{1}=4\lambda_{max}(C^{T}C)$ depend on $\sigma$. In Table 2, we show the effect $\sigma$ varying from 00 to $0.05$, causing the condition number $L_{1}/\mu$ to grow moderately. As a result, we see $T_{0}$ grow linearly in $L_{1}/\mu$ and $C_{0}$ grows exponentially, exceeding $10^{23}$.
For such problems, our theory predicts the subgradient method with $\alpha_{k}=2/\mu(k+2)$ may diverge in the first $T_{0}$ iterations but should eventually converge at least a $O(1/T)$ rate. Figure 2 numerically confirms this prediction with every performance measure exponentially growing to at least $10^{16}$ as $\sigma$ grows and a decreasing trend beginning before iteration $T_{0}$. We see $\bar{p}_{k}$, $\delta_{k}$ and $R_{k}^{2}(x_{\mathtt{OPT}})$ rapidly converge after $T_{0}$, whereas $p_{k}$ and $d_{k}$ only decrease sublinearly. This slow convergence is likely due to $p_{k}$ and $d_{k}$ being defined as weighted averages, which must slowly dilute the effects of early “bad” iterations. Our theory predicts such exponential divergences can be avoided by ensuring $T_{0}$ (and hence $C_{0}$) are small. For example, setting $\alpha_{0}=1/\mu$ and then $\alpha_{k}=\min\{1/L_{1},2/\mu(k+2)\}$ thereafter rather than $\alpha_{k}=2/\mu(k+2)$ as above ensures $T_{0}=0$. Figure 3 verifies this mitigates the previous diverging behavior.
@@
Acknowledgements.
This manuscript’s presentation and clarity were greatly improved by the careful thoughts and advice of two anonymous referees to whom the authors are very grateful.
Appendix A Deferred Proof of Lemma 2.4
Consider either $y\in\{x_{\mathtt{OPT}},x_{\mathtt{SL}}\}$. Suppose first $x_{k}$ is feasible and let $\bar{g}_{k}=\mathbb{E}_{\xi_{k}}g_{0}(x_{k};\xi_{k})+n_{y}\in\partial h_{y}(x_{k})$. Fix any $g_{y}\in\partial h_{y}(y)$. Note by the sum rule of subdifferential calculus, both $\bar{g}_{k}-\nabla f^{(2)}_{0}(x_{k})-n_{y}$ and $g_{y}-\nabla f^{(2)}_{0}(y)-n_{y}$ are subgradients of $f^{(1)}_{0}$ and hence both have norm bounded by $M$. Consider the $L$-smooth function
$\hat{h}_{y}(x)=f_{0}^{(2)}(x)+f^{(1)}_{0}(y)+\langle g_{y}-\nabla f_{0}^{(2)}(y),x-y\rangle{\color[rgb]{0.0,0.0,0.0}\definecolor[named]{pgfstrokecolor}{rgb}{0.0,0.0,0.0}\pgfsys@color@gray@stroke{0.0}\pgfsys@color@gray@fill{0.0}+r(y)}\ .$
Note since $f_{0}^{(1)}$ is convex with $g_{y}-\nabla f_{0}^{(2)}(y)-n_{y}\in\partial f^{(1)}_{0}(y)$, $h_{y}\geq\hat{h}_{y}$ and $h_{y}(y)=\hat{h}_{y}(y)$. Then one has
$\mathbb{E}_{\xi_{k}}\|g_{0}(x_{k};\xi_{k})+n_{y}\|_{2}^{2}$
$=\|\bar{g}_{k}\|_{2}^{2}+\mathbb{E}_{\xi_{k}}\|g_{0}(x_{k};\xi_{k}){\color[rgb]{0.0,0.0,0.0}\definecolor[named]{pgfstrokecolor}{rgb}{0.0,0.0,0.0}\pgfsys@color@gray@stroke{0.0}\pgfsys@color@gray@fill{0.0}+n_{y}}-\bar{g}_{k}\|_{2}^{2}$
$\leq 3\|\nabla\hat{h}_{y}(x_{k})\|_{2}^{2}+3\|\bar{g}_{k}-\nabla f^{(2)}_{0}(x_{k})-n_{y}\|^{2}+3\|g_{y}-\nabla f^{(2)}_{0}(y)-n_{y}\|_{2}^{2}+\sigma^{2}$
$\leq 6L(\hat{h}_{y}(x_{k})-\inf\hat{h}_{y})+6M^{2}+\sigma^{2}$
$\leq 6L\delta_{k}(y)+6L(h_{y}(y)-\inf\hat{h}_{y})+6M^{2}+\sigma^{2}$
where the first inequality bounds $\|a+b+c\|^{2}_{2}$ by $3\|a\|^{2}_{2}+3\|b\|^{2}_{2}+3\|c\|^{2}_{2}$ and uses the assumed variance bound, the second inequality uses smoothness to bound the first term as $\hat{h}_{y}(x_{k})-\frac{1}{2L}\|\nabla\hat{h}_{y}(x_{k})\|^{2}\geq\inf\hat{h}_{y}$ and the $M$-Lipschitzness of $f^{(1)}_{0}$ to bound the second and third terms, and the final inequality adds and subtracts ${\color[rgb]{0.0,0.0,0.0}\definecolor[named]{pgfstrokecolor}{rgb}{0.0,0.0,0.0}\pgfsys@color@gray@stroke{0.0}\pgfsys@color@gray@fill{0.0}6L}h_{y}(y)$ and upper bounds $\hat{h}_{y}(x_{k})$ by $h_{y}(x_{k})$.
Similarly, now suppose $x_{k}$ is infeasible and let $\bar{g}_{k}=\mathbb{E}_{\xi_{k}}g_{s(x_{k})}(x_{k};\xi_{k})\in\partial f_{s(x_{k})}(x_{k})$. Fix any $g_{y}\in\partial f_{s(x_{k})}(y)$. Note by the sum rule of subdifferential calculus, both $\bar{g}_{k}-\nabla f^{(2)}_{s(x_{k})}(x_{k})$ and $g_{y}-\nabla f^{(2)}_{s(x_{k})}(y)$ are subgradients of $f^{(1)}_{s(x_{k})}$ and hence both have norm bounded by $M$. Consider the $L$-smooth function
$\hat{f}_{s(x_{k})}(x)=f_{s(x_{k})}^{(2)}(x)+f^{(1)}_{s(x_{k})}(y)+\langle g_{y}-\nabla f_{s(x_{k})}^{(2)}(y),x-y\rangle\ .$
Note since $g_{y}-\nabla f_{s(x_{k})}^{(2)}(y)\in\partial f^{(1)}_{s(x_{k})}(y)$, $f_{s(x_{k})}\geq\hat{f}_{s(x_{k})}$ and $f_{s(x_{k})}(y)=\hat{f}_{s(x_{k})}(y)$. Then, identical reasoning to that above gives
$\mathbb{E}_{\xi_{k}}\|g_{0}(x_{k};\xi_{k})\|_{2}^{2}$
$=\|\bar{g}_{k}\|_{2}^{2}+\mathbb{E}_{\xi_{k}}\|g_{s(x_{k})}(x_{k};\xi_{k})-\bar{g}_{k}\|_{2}^{2}$
$\leq 3\|\nabla\hat{f}_{s(x_{k})}(x_{k})\|_{2}^{2}+3\|\bar{g}_{k}-\nabla f^{(2)}_{s(x_{k})}(x_{k})\|^{2}+3\|g_{y}-\nabla f^{(2)}_{s(x_{k})}(y)\|_{2}^{2}+\sigma^{2}$
$\leq 6L(\hat{f}_{s(x_{k})}(x_{k})-\inf\hat{f}_{s(x_{k})})+6M^{2}+\sigma^{2}$
$\leq 6L\delta_{k}(y)+6L(f_{s(x_{k})}(y)-\inf\hat{f}_{s(x_{k})})+6M^{2}+\sigma^{2}\ .$
Hence, Assumption C holds with
$L_{0}^{2}$ $=6M^{2}+\sigma^{2}+6L\max_{y\in\{x_{\mathtt{OPT}},x_{\mathtt{SL}}\}}\max_{s=1\dots m}\left\{h_{y}(y)-\inf\hat{h}_{y},f_{s}(y)-\inf\hat{f}_{s}\right\}$
$L_{1}$ $=6L\ .$